# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ 5 Empirical application
We implement our composite likelihood procedure on high-frequency data from the cryptocurrency market. [^13] These markets are ideal to investigate the pathwise properties of spot volatility, since trading is more or less never interrupted. In contrast, other asset classes feature periodic market closure. This can distort estimation of the persistence, since intermittent data are missing, potentially leading to incorrect matching of observations for the calculation of the sample acf. We look at the five largest free-floating coins in terms of market value at the end of 2023 (see, e.g., coingecko.com): Bitcoin (BTC), Ethereum (ETH), Binance (BNB), Solana (SOL), and Ripple (XRP). We examine the evolution of their spot exchange rate against Tether (USDT). The latter is a so-called “stablecoin”, whose value is pegged at parity against the US dollar, i.e. 1USDT = 1USD. [^14] According to Tether Limited, the token is backed one-to-one against fiat currency or cash equivalents (e.g., short-dated US Treasury bonds), although the alleged reserves were never audited. This is cause of much controversy and has lead to occasional sell-offs in USDT over time, where its price temporarily dropped far below parity against the USD. By and large, however, Tether maintains a fairly constant price at the official peg with limited volatility. We downloaded millisecond time-stamped tick-by-tick data free of charge from the Binance archive. [^15] https://data.binance.vision/?prefix=. Following Hansen, Kim, and Kimbrough (2022), we restrict attention to the period after January 1, 2019, where trading volume on the Binance platform was adequate. We collect data until December 31, 2023, so the sample consists of $T=1{,}826$ days. The exception is SOLUDST that is only available from August 11, 2020 and therefore has a slightly shorter span of $T=1{,}238$ days.
@@
BTCUSDT 1,781,576 48.5 15.8 $\underset{(0.296)}{0.323}$ $\underset{(0.010)}{1.265}$ $\underset{(0.009)}{-0.384}$ $\underset{(0.074)}{0.337}$ $\underset{(0.060)}{1.256}$ $\underset{(0.012)}{-0.218}$
ETHUSDT 663,552 130.2 19.8 $\underset{(0.281)}{0.243}$ $\underset{(0.010)}{1.225}$ $\underset{(0.010)}{-0.383}$ $\underset{(0.072)}{0.306}$ $\underset{(0.062)}{1.216}$ $\underset{(0.011)}{-0.220}$
BNBUSDT 372,319 232.1 21.6 $\underset{(0.282)}{0.204}$ $\underset{(0.009)}{1.095}$ $\underset{(0.010)}{-0.385}$ $\underset{(0.069)}{0.287}$ $\underset{(0.051)}{1.088}$ $\underset{(0.012)}{-0.226}$
SOLUSDT 348,259 248.1 29.7 $\underset{(0.252)}{0.092}$ $\underset{(0.011)}{1.300}$ $\underset{(0.011)}{-0.366}$ $\underset{(0.070)}{0.187}$ $\underset{(0.077)}{1.294}$ $\underset{(0.013)}{-0.209}$
XRPUSDT 309,173 279.5 24.6 $\underset{(0.322)}{0.459}$ $\underset{(0.009)}{1.167}$ $\underset{(0.009)}{-0.378}$ $\underset{(0.075)}{0.368}$ $\underset{(0.054)}{1.156}$ $\underset{(0.012)}{-0.205}$
@@
In the left-hand side of Table 5, we provide a summary of the amount of transaction data available. The selected pairs are vastly liquid. For example, BTCUSDT averages 1,781,576 transactions per day (median of 1,045,734), equivalent to an intertrade duration of 48.5 milliseconds (median of 82.6). Meanwhile, the least liquid pair XRPUSDT still has about four trades per second. [^16] As a comparison, from January 1, 2019 to October 14, 2021 the average daily number of transactions in the front contract of the E-mini S&P 500 futures (ES) was 318,157 (median of 267,100 and maximum of 1,604,309). Meanwhile, the dollar volume in the ES was about 2.5 times larger than in the BTCUSDT.
Even though the bid-ask spread in the selected coins is routinely less than a basis point, sampling the spot exchange rate process at the tick-by-tick level is evidently going to instill a nontrivial amount of microstructure noise in the derived realized variance measure (see, e.g., Hansen and Lunde, 2006). To avoid this, we create a 15-second equidistant transaction price series, corresponding to $n=5{,}760$ daily log-price increments, by pre-averaging observations that are nearest to each time point (Jacod, Li, Mykland, Podolskij, and Vetter, 2009; Mykland and Zhang, 2016; Podolskij and Vetter, 2009a, b).
Figure 1: Properties of BTCUSDT spot variance.
Panel A: Diurnal variation. Panel B: Log-daily realized variance.
@@
Note. In Panel A, we plot a nonparametric estimator of the intraday periodicity in the spot variance of BTCUSDT. We sample the price process every 15th second from January 1, 2019 to December 31, 2023. Time 0 is midnight in Coordinated Universal Time (UTC). The line is a 5-minute moving average. In Panel B, we plot the daily realized variance (expressed in annualized standard deviation terms) computed after removal of the estimated diurnal pattern. The latter has been log-transformed.
@@
In the Panel A of Figure 1, we plot an estimate of the intraday periodicity in the spot variance of the cryptocurrency market. Here, BTCSUDT serves as a illustration, but in general the volatility of the various exchange rates exhibit near-identical behavior, indicative of a strong commonality. We employ the nonparametric estimator of Taylor and Xu (1997), which averages the time-of-the-day squared 15-second log-returns (a proxy for the average point-in-time intraday variance) and normalizes the sum of these numbers to one to form a seasonality estimate. As readily seen, there is a discernible nonstationary component in the intraday evolution of the volatility of BTCUSDT, which features a complex intraday seasonality pattern. This clashes with our assumed stationarity in (2.1), which can distort the estimation of the stationary Gaussian processes. To avoid this, we pre-filter the 15-second log-return series with the estimated diurnal coefficient to ensure that the rescaled time series is closer to stationary.
@@
We next construct a non-overlapping 2-hour realized variance, which is input into the numerical optimization procedure. To ensure that price jumps do not interfere with our estimation results, we zero out too large absolute log-returns with the truncation approach of Mancini (2009). In Panel B of Figure 1, we plot the resulting log-daily realized variance series, adding up the local diurnal-corrected realized variance estimates to the daily horizon. As expected, it displays a lot of persistence. The typical value of realized variance across currencies is shown in Table 5. The estimates imply that the diffusive variance is rather high, on average, and somewhat in line with the overall level of the total quadratic return variation often seen in individual stocks.
@@
The two stationary Gaussian processes from Section 3 are estimated following the design from Section 4. To get a gauge at the magnitude of the standard errors inherent in the parameter estimates, we employ a parametric bootstrap with 1,000 repetitions. The results are presented in the right-hand side of Table 5. We only include the parameters relating to the acf, since the mean is always estimated in line with the sample average of log-realized variance.
@@
We observe that the roughness index of log-spot variance is located around -0.40 for the fOU process, which we recall is related to the Hurst exponent via the relation $\alpha=H-1/2$. Hence, as consistent with the recent literature on rough volatility, we uncover that there is overwhelming evidence in support of the variance process being more erratic than a standard Brownian motion at short time scales (see, e.g. Bolko, Christensen, Pakkanen, and Veliyev, 2023; Fukasawa, Takabatake, and Westphal, 2022; Gatheral, Jaisson, and Rosenbaum, 2018; Wang, Xiao, and Yu, 2023). The $\kappa$ parameter is very close to zero, so the process exhibits near-unit root behavior.
@@
Turning next to the Cauchy class, the results are striking. On the one hand, $\alpha$ is now estimated around -0.20. While the values are much higher than for the fOU process, they are still located far in roughness space. On the other hand, $\beta$ is estimated far in the long memory range around 0.30 – 0.35. Noting that the link between $\beta$ and the fractional differencing parameter $d$ in a discrete-time ARFIMA($p$,$d$,$q$) process is $\beta=1-2d$ (see Brockwell and Davis, 1991, equation 13.2.1), this corresponds to a $d$ of about 0.30 – 0.40, on average. This agrees with early evidence from the literature on modeling and forecasting realized variance as a fractionally integrated process, see, e.g., Andersen, Bollerslev, Diebold, and Labys (2000), who report that $d\simeq 0.40$ for foreign exchange rates.
Figure 2: Heatmap of composite likelihood function for BTCUSDT.
@@
Note. We create a heatmap of the composite likelihoood function for the fOU process (in Panel A) and Cauchy class (in Panel B). A cooler/warmer temperature (blue/red) indicates that the likelihood is lower/higher. We fix the mean and standard deviation ($\mu$ and $\nu$) at the sample average and sample standard deviation of the log-realized variance and vary the remaining free parameters over a broad range of the parameter space. The MCLE is reported with a black cross.
@@
In Figure 2, we construct a heatmap of the composite likelihood function for the BTCUSDT spot exchange rate. Panel A is for the fOU process, while Panel B is for the Cauchy class. To reduce the plot to a two-dimensional plane, we fix $\mu$ and $\nu$ at the sample average and sample standard deviation of the log-realized variance, while varying $\kappa$ and $\beta$ over a broad range of the parameter space and $\alpha$ through the permissible area. The color code indicates the height of the likelihood function with higher (lower) contour values being indicated as a warmer (cooler) temperature. [^17] We should emphasize that this piece of analysis is not readily available for the MME approach.
@@
As consistent with the parameter estimates of BTCUSDT from Table 5, Panel A of Figure 2 indicates that the CL of the fOU process prefers a rough dynamic with a Hurst exponent less than a half. The optimal value of the mean reversion hovers close to the boundary of the parameter space at zero, where the fOU process becomes a scaled version of the non-stationary fBm. Shi and Yu (2022) look at the discrete-time ARFIMA($1$,$d$,00) model: $(1-\phi L)(Y_{t}-\mu)=\sigma_{\epsilon}(1-L)^{-d}\epsilon_{t}$, where $L$ is the lag operator and $\epsilon_{t}\sim D(0,1)$ is a white noise. They note that $\phi=0$ and $d=0.5$ is observationally equivalent to $\phi=1$ and $d=-0.5$, leading to identification failure, see also Liu, Shi, and Yu (2020) and Li, Phillips, Shi, and Yu (2023). This can be interpreted as either fast mean reversion and long memory or no mean reversion and anti-persistence. This agrees with a first impression of the likelihood function: the fOU model is prepared to trade roughness for additional mean reversion with only a minimal drop in likelihood and the surface appears to exhibit a plateau over an extended region in that direction. This is a symptom that the model is scrambling to fit both ends of the acf.
@@
However, to get a better gauge at the exact curvature, in Figure 3 we include the profile composite likelihood function of the fOU process in Panel A, $CL_{\alpha}(\kappa)=\max_{\alpha}CL(\alpha,\kappa)$ and the Cauchy class in Panel B, $CL_{\alpha}(\beta)=\max_{\alpha}CL(\alpha,\beta)$. We again set $\mu$ and $\nu$ equal to the sample average and sample standard deviation of the log-realized variance and suppress their influence on the composite likelihood. We then plot this as a function of $\kappa$ or $\beta$ after maximizing with respect to $\alpha$, where the optimal value of the latter is shown on the right-hand $y$-axis. To make the graphs comparable, we normalize the profile likelihood to reach a maximum of one. Overall, the evolution of the fOU profile likelihood in Panel A shows that it is, in fact, monotonically increasing toward the boundary of the parameter space, whereas the Cauchy class has a maximum in the interior. It is interesting to observe that as $\kappa$ increases for the fOU, the optimal value of $\alpha$ does not venture into the long memory region but levels out around zero.
Figure 3: Profile composite likelihood function for BTCUSDT.
@@
Note. We show the profile composite likelihood function of the fOU process in Panel A, $CL_{\alpha}(\kappa)=\max_{\alpha}CL(\alpha,\kappa)$ and the Cauchy class in Panel B, $CL_{\alpha}(\beta)=\max_{\alpha}CL(\alpha,\beta)$. We fix $\mu$ and $\nu$ at the sample average and sample standard deviation of the log-realized variance. We then plot the composite likelihood as a function of $\kappa$ or $\beta$ after maximizing with respect to $\alpha$, where the optimal value of the latter is shown on the right-hand $y$-axis.
At the end of the day, the fOU process has to make a decision: roughness or long memory. It selects roughness to capture the initial decline in the acf and to fit the subsequent slow decay it forces the mean reversion parameter near zero to approximate long-range dependence. However, the Cauchy class does not suffer from this shortcoming and is able to separate these effects to demonstrate that both are required to describe the data. The $(\alpha,\beta)$ estimate is within the interior of the parameter space and appears to be identified with a high degree of accuracy, gauging from the heatmap, profile likelihood, and the magnitude of the standard errors.

### v2

@@ 5 Empirical application
We implement our composite likelihood procedure on high-frequency data from the cryptocurrency market. [^24] These markets are ideal to investigate the (pathwise) properties of a stochastic process, since trading is more or less never interrupted. In contrast, other asset classes feature periodic market closure. This can distort estimation of the persistence, since intermittent data are missing, potentially leading to incorrect matching of observations for the calculation of the sample ACF. We look at the five largest free-floating coins in terms of market value at the end of 2024 (see, e.g., coingecko.com): Bitcoin (BTC), Ethereum (ETH), Solana (SOL), Ripple (XRP), and Binance (BNB). We examine the evolution of their spot exchange rate against Tether (USDT). The latter is a so-called “stablecoin”, whose value is pegged at parity against the US dollar, i.e. 1USDT = 1USD. [^25] According to Tether Limited, the token is backed one-to-one against fiat currency or cash equivalents (e.g., short-dated US Treasury bonds), although the alleged reserves were never audited. This is the cause of much controversy and has lead to occasional sell-offs in USDT, where its price temporarily dropped far below parity against the USD. By and large, however, Tether maintains a fairly constant price at the official peg with limited volatility. We downloaded millisecond time-stamped tick-by-tick data free of charge from the Binance archive. [^26] https://data.binance.vision/. Following Hansen, Kim, and Kimbrough (2022), we restrict attention to the period after January 1, 2019, where trading volume on the Binance platform was adequate. We collect data until March 27, 2025, so the sample consists of T=2,278T=2{,}278 days. The exception is SOLUDST that is only available from August 11, 2020 and therefore has a slightly shorter span of T=1,690T=1{,}690 days.
@@
$100\times\kappa$ $\nu$ $\alpha$ $\beta$ $\nu$ $\alpha$ $100\times\kappa$ $\nu$ $\alpha$ $\beta$ $\nu$ $\alpha$
BTC 2,044,826 42.3 52.0 174.2 $\underset{(0.050)}{0.652}$ $\underset{(0.008)}{1.182}$ $\underset{(0.001)}{-0.375}$ $\underset{(0.009)}{0.236}$ $\underset{(0.017)}{1.181}$ $\underset{(0.002)}{-0.293}$ $\underset{(0.000)}{0.010}$ $\underset{(0.006)}{0.972}$ $\underset{(0.001)}{-0.415}$ $\underset{(0.003)}{0.103}$ $\underset{(0.009)}{0.972}$ $\underset{(0.001)}{-0.364}$
ETH 977,713 88.4 65.2 80.6 $\underset{(0.060)}{0.449}$ $\underset{(0.008)}{1.143}$ $\underset{(0.002)}{-0.377}$ $\underset{(0.008)}{0.206}$ $\underset{(0.016)}{1.142}$ $\underset{(0.001)}{-0.300}$ $\underset{(0.000)}{0.003}$ $\underset{(0.009)}{1.119}$ $\underset{(0.001)}{-0.417}$ $\underset{(0.002)}{0.083}$ $\underset{(0.013)}{1.119}$ $\underset{(0.001)}{-0.371}$
SOL 693,803 124.5 110.7 29.8 $\underset{(0.060)}{1.079}$ $\underset{(0.006)}{1.016}$ $\underset{(0.001)}{-0.345}$ $\underset{(0.008)}{0.241}$ $\underset{(0.013)}{1.015}$ $\underset{(0.001)}{-0.256}$ $\underset{(0.001)}{0.020}$ $\underset{(0.015)}{1.353}$ $\underset{(0.003)}{-0.386}$ $\underset{(0.003)}{0.086}$ $\underset{(0.023)}{1.353}$ $\underset{(0.001)}{-0.329}$
XRP 466,774 185.1 82.9 24.1 $\underset{(0.050)}{1.167}$ $\underset{(0.007)}{1.154}$ $\underset{(0.001)}{-0.355}$ $\underset{(0.010)}{0.262}$ $\underset{(0.016)}{1.153}$ $\underset{(0.002)}{-0.266}$ $\underset{(0.004)}{0.060}$ $\underset{(0.008)}{1.153}$ $\underset{(0.002)}{-0.385}$ $\underset{(0.004)}{0.116}$ $\underset{(0.016)}{1.152}$ $\underset{(0.001)}{-0.323}$
BNB 438,524 197.0 69.8 22.8 $\underset{(0.060)}{0.630}$ $\underset{(0.007)}{1.093}$ $\underset{(0.001)}{-0.362}$ $\underset{(0.007)}{0.214}$ $\underset{(0.014)}{1.092}$ $\underset{(0.001)}{-0.279}$ $\underset{(0.000)}{0.007}$ $\underset{(0.009)}{1.149}$ $\underset{(0.001)}{-0.409}$ $\underset{(0.003)}{0.088}$ $\underset{(0.014)}{1.148}$ $\underset{(0.001)}{-0.358}$
@@
In the left-hand side of Table 5, we provide a summary of the amount of transaction data available. The selected pairs are vastly liquid. For example, BTC averages 2,044,826 transactions per day (median of 1,257,976), equivalent to an intertrade duration of 42.3 milliseconds (median of 68.7). Meanwhile, the least liquid pair BNB still has more than five trades per second. [^27] As a comparison, from January 1, 2019 to October 14, 2021 the average daily number of transactions in the front contract of the E-mini S&P 500 futures (ES) was 318,157 (median of 267,100 and maximum of 1,604,309). The dollar volume in the ES was about 2.5 times larger than in the BTC.
We follow Wang, Xiao, and Yu (2023); Shi, Yu, and Zhang (2024) and apply our composite likelihood framework to both the volatility and trading volume of these cryptocurrencies. While the dynamic of volatility is of primary interest, in practice the spot variance is latent, and it needs to be inferred from the discretely observed high-frequency data. This may distort the subsequent analysis, as further discussed below. By contrast, the trading volume is immediately accessible, and it is known to be strongly correlated with volatility. The evidence from the latter further buttresses our findings for the estimated spot variance.
5.1 Analysis of spot variance
In this subsection, we follow the simulation section and let $Y$ be the log-variance of the log-spot exchange rate process, meaning that $Y_{t}=\log\sigma_{t}^{2}$. In practice, this variable is unobserved, and to recover it, we employ a standard RV calculated over short non-overlapping time intervals, see, e.g., Andersen and Bollerslev (1998) and Barndorff-Nielsen and Shephard (2002).
Suppose there are $m$ equidistant increments of a log-price in the interval $t,t+\Delta]$. Then, the RV over this interval is defined as
$RV_{t,t+\Delta}^{m}=\sum_{j=1}^{m}(\log P_{t+\frac{j}{m}\Delta}-\log P_{t+\frac{j-1}{m}\Delta})^{2},$ (5.1)
for $t=\Delta,2\Delta,\dots,n\Delta$.
In theory, RV is able to estimate the latent spot, or point-in-time, variance. This follows from a double-asymptotic framework in which an increasing number of log-price increments are observed over a shrinking time interval, i.e. $\Delta\rightarrow 0$ and $m\rightarrow\infty$ such that $\Delta m\rightarrow\infty$. In this case, under weak regularity conditions, $RV_{t,t+\Delta}^{m}\overset{\mathbb{P}}{\longrightarrow}\sigma_{t}^{2}$ (see Jacod and Protter, 2012). By the continuous mapping theorem, $\log RV_{t,t+\Delta}^{m}\overset{\mathbb{P}}{\longrightarrow}\log\sigma_{t}^{2}=Y_{t}$. In fact, RV is the MLE in a parametric model with no drift and constant volatility, so by the invariance principle, the logarithm of RV is the MLE of the log-spot variance. In this sense, RV is the best possible proxy available.
In practice, however, RV contains at least two important sources of error. First, $\Delta$ is normally fixed and RV is instead consistent for the integrated variance (IV), $RV_{t,t+\Delta}^{m}\overset{\mathbb{P}}{\longrightarrow}IV_{t,t+\Delta}=\int_{t}^{t+\Delta}\sigma_{s}^{2}\mathrm{d}s$ as $m\rightarrow\infty$. The integration introduces a smoothing effect, which tends to bias estimators of the Hurst parameter upward. Second, with real data it is ill-advised to sample too often, because of microstructure noise (Hansen and Lunde, 2006), which puts an upper bound on $m$. While RV converges in probability to IV in the infill asymptotic limit, $\Delta$ fixed and $m\rightarrow\infty$, with $m$ finite it contains an estimation error. Thus, the blockwise RV fluctuates around the true value of IV, which tends to bias estimators of the Hurst parameter downward. [^28] This has been dubbed elusive roughness in Fukasawa, Takabatake, and Westphal (2022). Although these effects may partially cancel out, they can severely impair the estimation.
Even though the bid-ask spread in the included coins is routinely less than a basis point, the presence of other market frictions, such as price discreteness and bid-ask bounce, can evidently instill a nontrivial bias in RV if we insist on sampling the spot exchange rate process at the maximum available tick-by-tick level in order to mitigate the above concerns. To circumvent this problem and strike a balance, we construct a 15-second equidistant log-price series, corresponding to 5,760 daily log-price increments. This choice was guided by a volatility signature plot, which is a diagnostic tool designed to gauge the impact of microstructure noise on RV (see, e.g., Hansen and Lunde, 2006). It computes the sample average daily RV as a function of the sampling frequency:
$\overline{RV}(m)=T^{-1}\sum_{t=1}^{T}RV_{t-1,t}^{m},$ (5.2)
where $RV_{t-1,t}^{m}$ is the RV over the whole day $t$ sampled with a time gap $1/m$.
In general, the behavior of a volatility signature plot depends on the properties of the noise. If it is absent (or negligible), $\overline{RV}(m)$ is an estimator of the average IV, so the graph of the function should be a flat line up to sampling variation. A noise that is independent of the efficient price causes an upward bias in the estimate and, if the noise is itself i.i.d., $\overline{RV}(m)\overset{\mathbb{P}}{\longrightarrow}\infty$ as $m\rightarrow\infty$.
Figure 2: Microstructure noise and intraday variation in volatility.
Panel A: Volatility signature. Panel B: Diurnal variation.
@@ 5.1 Analysis of spot variance
Note. In Panel A, we present a volatility signature plot of BTC, i.e. the sample average of daily RV as a function of the sampling frequency ranging from 10-minute (slowest) to one-second (fastest). Each observation is scaled by the 10-minute value. We show a two standard error band around the point estimate. In Panel B, we plot a nonparametric estimator of the intraday periodicity in the associated 15-second squared log-return process. The line is a 5-minute moving average. The sample runs from January 1, 2019 to March 27, 2025. Time 0 is midnight in Coordinated Universal Time (UTC).
In Panel A of Figure 2, we present the volatility signature plot of BTC, which is the coin used throughout as an illustration. The charts for the other cryptocurrencies are contained in Appendix D. We vary $m$ from 144 to 86,400, which correspond to 10-minute sampling as the slowest to 1-second sampling as the fastest. Each $\overline{RV}(m)$ has been scaled by the 10-minute estimate such that, as $m$ increases, the evolution of the graph can be interpreted as the relative bias induced by the accumulation of microstructure noise. To get a crude impression of the sampling variation, we surround each point estimate by a two standard error confidence band. Overall, there is little evidence to suggest that microstructure noise wreaks havoc on the RV of BTC with 15-second—or even much faster—sampling of the price process. This impression is corroborated by the remaining figures in the appendix. The exception is SOL, where we do notice an uptick in $\overline{RV}(m)$ at the 15-second horizon, and the estimate is slightly outside the 10-minute confidence interval, so we stop there. [^29] As an extra robustness check, we computed the Hausman test for the presence of microstructure noise in high-frequency data developed by Aït-Sahalia and Xiu (2019). A caveat with their theoretical framework is that the volatility process is assumed to be an Itô semimartingale, so this piece of analysis should be taken with a grain of salt. Nonetheless, when evaluating their stochastic volatility- and price jump-robust test statistic at the 1% significance level, the null hypothesis of no microstructure noise is rejected only 2.19% of the days in the BTC sample at the 15-second frequency. This is rather close to the nominal level, especially since their test statistic was found to be slightly oversized in finite samples.
To remove traces of noise even at the 15-second horizon, we follow Jacod, Li, Mykland, Podolskij, and Vetter (2009) by pre-averaging transaction prices that are closest to each time point on the sampling grid.
@@
In the Panel B of Figure 2, we plot an estimate of the intraday periodicity in the second moment of the BTC log-return series. [^30] In general, the intraday volatility of the various exchange rates exhibit near-identical behavior, indicative of a strong commonality. We employ a nonparametric estimator that averages the time-of-the-day squared 15-second log-returns (a proxy for the average point-in-time intraday variance) and normalizes the sum of these numbers to one to form a seasonality estimate. As readily seen, there is a discernible nonstationary component in the intraday evolution of the volatility of BTC, which features a complex behavior. This clashes with our assumed stationarity in (2.1), which can distort the estimation of the stationary Gaussian processes. To avoid this, we pre-filter the 15-second log-return series with the estimated diurnal coefficient to ensure that the rescaled time series is closer to stationary. In addition to this, we observe a marked day-of-the-week effect in the variance process, which we too eliminate.
@@
We next construct a non-overlapping two-hour RV for the corrected 15-second log-returns. [^31] To ensure that price jumps do not interfere with our estimation results, we also zero out too large absolute log-returns with the truncation approach of Mancini (2009). Although the block size is a bit arbitrary, it is an attempt to balance the inherent errors in the RV measure, as discussed above. The concrete choice follows extant literature on spot volatility estimation (see, e.g., Christensen, Thyrsgaard, and Veliyev, 2019). The typical value of RV across currencies at this sampling frequency is shown in Table 5. The estimates imply that the diffusive variance is rather high on average and exceeds the overall level of the total quadratic return variation often seen in even very risky individual stocks.
Figure 3: Distribution and persistence of log-RV.
Panel A: Kernel density estimate. Panel B: Sample ACF.
Refer to caption Refer to caption
Note. In Panel A, we present a kernel density estimate of the two-hour log-RV. A standard normal distribution is superimposed as a benchmark. In Panel B, we plot the sample ACF of this measure. “Corrected” is the version that is purged from intraday periodicity and a day-of-the-week effect in volatility. As a comparison, we also show the ACF of the raw “Uncorrected” RV.
How close to Gaussian is the two-hour log-RV? To get a crude impression, we center and scale the time series by its sample average and sample standard deviation. In Panel A of Figure 3, we plot the histogram of the standardized statistic. The distribution of spot log-RV is very close to the bell-shaped curve of the Gaussian density function, which is consistent with previous literature (e.g. Andersen, Bollerslev, Diebold, and Ebens, 2001; Andersen, Bollerslev, Diebold, and Labys, 2003).
We next examine the correlation structure of the process. In Panel B of Figure 3, we present a correlogram of the spot log-RV. The lag length goes up to a full year. Here, “Uncorrected” represents the raw estimator, while “Corrected” is based on the version that accounts for intraday periodicity and the day-of-the-week effect. Both point toward a strong persistence in the extracted log-spot variance process, which—as noted above—tends to understate the population autocorrelation in the presence of long memory. This effect is arguably excerbated here, because we are working with a noisy proxy of the volatility (e.g., Hansen and Lunde, 2014). It is evidently hard to say what type of model is “best” based on the graph alone, but the ability of being able to disentangle the short- and long-end of the memory lane does appear useful in order to fit the initial steep decline of the curve combined with the slow leveling off.
@@
The two stationary Gaussian processes from Section 3 are estimated following the design from Section 4. We input the corrected log-RV into the numerical optimization procedure. To get a gauge at the magnitude of the standard errors inherent in the parameter estimates, we use the limiting covariance expressions from Theorem 2.2. Here, we plug our estimates into the limiting covariance expression, where we numerically compute the derivatives of the autocovariance terms. The results are presented in the right-hand side of Table 5. We only include the parameters relating to the ACF, since the mean is always estimated in line with the sample average of log-RV (with a weighted average of log-RV, as explained in Footnote 22).
@@
We observe that the roughness index of spot log-RV is located around 0.35-0.35 for the fOU process, which we recall is related to the Hurst exponent via the relation $\alpha=H-1/2$. Hence, as consistent with the recent literature on rough volatility, we uncover that there is overwhelming evidence in support of the variance process being more erratic than a sBm at short time scales (see, e.g. Bolko, Christensen, Pakkanen, and Veliyev, 2023; Fukasawa, Takabatake, and Westphal, 2022; Gatheral, Jaisson, and Rosenbaum, 2018; Wang, Xiao, and Yu, 2023). The $\kappa$ parameter is very close to zero, so the process exhibits near-unit root behavior.
@@
Turning next to the Cauchy class, the results are striking. On the one hand, $\alpha$ is now estimated around 0.27-0.27. While the values are higher than for the fOU process, they are still located far in roughness space. On the other hand, $\beta$ is estimated far in the long memory range around 0.20 – 0.25. Noting that the link between $\beta$ and the fractional differencing parameter $d$ in a discrete-time ARFIMA($p$,$d$,$q$) process is $d=(1-\beta)/2$ (see Brockwell and Davis, 1991, equation 13.2.1), this corresponds to a $d$ close to 0.40 on average. This agrees with early evidence from the literature on modeling and forecasting log-RV as a fractionally integrated process, see, e.g., Andersen, Bollerslev, Diebold, and Ebens (2001); Andersen, Bollerslev, Diebold, and Labys (2003).
Figure 4: Heatmap of composite likelihood function.
@@
Note. We create a heatmap of the composite likelihoood function for the fOU process (in Panel A) and Cauchy class (in Panel B). A cooler/warmer temperature (blue/red) indicates that the likelihood is lower/higher. We fix the mean and standard deviation ($\mu$ and $\nu$) at the sample average and sample standard deviation of the log-RV and vary the remaining free parameters over a broad range of the parameter space. The MCLE is reported with a black cross.
@@
In Figure 4, we construct a heatmap of the composite likelihood function for the log-RV of the BTC spot exchange rate. Panel A is for the fOU process, while Panel B is for the Cauchy class. To reduce the plot to a two-dimensional plane, we fix $\mu$ and $\nu$ at the sample average and sample standard deviation of the log-RV, while varying $\kappa$ and $\beta$ over a broad range of the parameter space and $\alpha$ through the permissible area. The color code indicates the height of the likelihood function with higher (lower) contour values being indicated as a warmer (cooler) temperature. [^32] We should emphasize that this piece of analysis is not readily available for the MME approach.
@@
As consistent with the parameter estimates of BTC from Table 5, Panel A of Figure 4 indicates that the composite likelihood of the fOU process prefers a rough dynamic with a Hurst exponent less than a half. The optimal value of the mean reversion hovers close to the boundary of the parameter space at zero, where the fOU process becomes a scaled version of the non-stationary fBm. Shi and Yu (2023) look at the discrete-time ARFIMA(11,dd,0) model: $(1-\phi L)(Y_{t}-\mu)=\sigma_{\epsilon}(1-L)^{-d}\epsilon_{t}$, where $L$ is the lag operator and $\epsilon_{t}\sim D(0,1)$ is a white noise. They note that $\phi=0$ and $d=0.5$ is observationally equivalent to $\phi=1$ and d=0.5d=-0.5, leading to identification failure, see also Liu, Shi, and Yu (2020) and Li, Phillips, Shi, and Yu (2025). This can be interpreted as either fast mean reversion and long memory or no mean reversion and anti-persistence. This agrees with a first impression of the likelihood function: the fOU model is prepared to trade roughness for additional mean reversion with only a minimal drop in likelihood and the surface appears to exhibit a plateau over an extended region in that direction. This is a symptom that the model is scrambling to fit both ends of the ACF.
@@
To get a gauge at the exact curvature, in Figure 5 we include the profile composite likelihood function of the fOU process in Panel A, $CL_{\alpha}(\kappa)=\max_{\alpha}CL(\alpha,\kappa)$ and the Cauchy class in Panel B, $CL_{\alpha}(\beta)=\max_{\alpha}CL(\alpha,\beta)$. We again set $\mu$ and $\nu$ equal to the sample average and sample standard deviation of the log-RV and suppress their influence on the composite likelihood. We then plot this as a function of $\kappa$ or $\beta$ after maximizing with respect to $\alpha$, where the optimal value of the latter is shown on the right-hand $y$-axis. To make the graphs comparable, we normalize the profile likelihood to reach a maximum of one. Overall, the evolution of the fOU profile likelihood in Panel A shows that it is, in fact, monotonically increasing toward the boundary of the parameter space, whereas the Cauchy class has a maximum in the interior. It is interesting to observe that as $\kappa$ increases for the fOU, the optimal value of $\alpha$ does not venture into the long memory region but levels out around zero, as implied by a sBm.
Figure 5: Profile composite likelihood function.
@@
Note. We show the profile composite likelihood function of the fOU process in Panel A, $CL_{\alpha}(\kappa)=\max_{\alpha}CL(\alpha,\kappa)$ and the Cauchy class in Panel B, $CL_{\alpha}(\beta)=\max_{\alpha}CL(\alpha,\beta)$. We fix $\mu$ and $\nu$ at the sample average and sample standard deviation of the log-RV. We then plot the composite likelihood as a function of $\kappa$ or $\beta$ after maximizing with respect to $\alpha$, where the optimal value of the latter is shown on the right-hand $y$-axis.
At the end of the day, the fOU process has to make a decision: either roughness or long memory. It selects roughness to capture the initial decline in the ACF and to fit the subsequent slow decay it forces the mean reversion parameter near zero to approximate long-range dependence. However, the Cauchy class does not suffer from this shortcoming and is able to separate these effects to demonstrate that both are required to describe the data. The $(\alpha,\beta)$ estimate is within the interior of the parameter space and appears to be identified with a high degree of accuracy, gauging from the heatmap, profile likelihood, and the magnitude of the standard errors.